Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
Mary’s hourly wage is twice that of John’s. John’s and Dennis’ hourly wages together total $60. If Dennis earns 1/3 of John’s rate, what does John make per hour?

Answer by stanbon(75887)

Mary’s hourly wage is twice that of John’s.
Let John's hourly wage be "x"; Then Mary's hourly wage is "2x"
----------------------
John’s and Dennis’ hourly wages together total $60. If Dennis earns 1/3 of John’s rate, what does John make per hour?
EQUATIONS:
J + D = 60
D = (1/3)J
------------
Substitute to solve for J:
J + (1/3)J = 60
(4/3)J = 60
J = (3/4)60 = $45 per hour (John's hourly wage rate)
Substitute in J+D=60 to solve for D:
45 + D = 60
D = $15 (Dennis's hour wage rate)
===================
Cheers,
Stan H.